OpenStudy (anonymous):

f(x) = -16x2 + 22x + 3 Part A: What are the x-intercepts of the graph of the f(x)? Show your work. (2 points) Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work. (3 points) Part C: What are the steps you would use to graph f(x)? Justify that you can use the answers obtained in Part A and Part B to draw the graph. (5 points)

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

which part?

OpenStudy (anonymous):

part b finding the vertex

OpenStudy (anonymous):

you know that this graph represents a parabola?

OpenStudy (anonymous):

yes but i dont know how to find the vertex without using a graph

OpenStudy (anonymous):

for the vertex of the parabola, it has to reach a maxima or a minima. can you tell me what to do to find out the point where the graph reaches its maxima or minima?

OpenStudy (anonymous):

no i donk know how to find that. ik that it is a minimum but thats it. is their a formula i can use

OpenStudy (anonymous):

i think it is a maxima, but let's get there later. you know how to differenciate right?

OpenStudy (anonymous):

i think maximum is downwards and has a positive leading coefficient and minimum is upwards and has a negative leading coefficient

OpenStudy (anonymous):

i will help you with that later. at first tell me do you know how to differenciate?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay, no problem. at first, let me be clear that this graph will attain maximum as it is downwards, because its co-efficient of leading term is negative

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

when i graph it it look like its going upwards

OpenStudy (anonymous):

now, for maximum, you need to find out where y=f(x)=-16 x^2+ 22x +3 attains maximum. -16x^2+22x+3=-(4x)^2+2*4x*2.75-(2.75)^2+3+(2.75)^2=10.5625-(4x-2.75)^2 with a little calculation you can reach this step.. tell me if you have understood

OpenStudy (anonymous):

ok i get it

OpenStudy (anonymous):

i just completed the square adjusted with 16x^2 and 22x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

nice! now as you can see, (4x-2.75)^2 is greater than or equal to zero always, so 10.5625-(4x-2.75)^2 will be maximum when 4x-2.75=0, or x=2.75/4 got it?

OpenStudy (anonymous):

yes i got it

OpenStudy (anonymous):

so, what is the maximum value of f(x)? obviously when 4x-2.75=0, that means x=0.6875 what is the value of f(x) at that point? clearly 10.5625 as you can see the function attains a maximum at (0.6875, 10.5625), it is the vertex

OpenStudy (anonymous):

ohh ok so basically you find the maximum point?

OpenStudy (anonymous):

yeah, maximum or minimum, depending on the co-efficient of the leading term. if it's positive, the graph is upward, f(x) is has minima, else the opposites

OpenStudy (anonymous):

oh ok thank you so much can u help with part c now

OpenStudy (anonymous):

this is what i put I would graph this equation by seeing what in formation i have. I already know the vertex and the points where it crosses the x axis. We also know that it is going to be an upwards parabola. I can plot all the points and make a rough drawing of how the parabola line would look.

OpenStudy (anonymous):

no, this is downward parabola

OpenStudy (anonymous):

oh ok is the rest good

OpenStudy (anonymous):

everything else you have said are all correct :) good job!

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

you are welcome :)